A collection of dimes and nickels is worth $1.70. There are 21 coins in all. How many each coin are there?
Please I really need your help because I will recite this tomorrow hoping for your help really thank you!

Answer by ikleyn(52864)

.

Let x be the number of dimes.

Then the number of nickels is (21-x) coins.


With it, we can write the total money equation


    10x + 5*(21-x) = 170 cents


(5 is standing for the dimes worth and 10 stands for the dimes worth).


From the equation,  



    10x + 105 - 5x = 170

    10x - 5x = 170 - 105

    5x       = 65

     x       = 65/5 = 13.


ANSWER.  13 dimes and the rest coins (21-13) = 8 are nickels.


CHECK.   13*10 + 8*5 = 170 centsa = $1.70.   ! Correct !

Solved.

There are other ways to solve the problem, but this one is one of the standard Algebra methods.

-------------

To see other way, look into the lessons
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive
    - Three methods for solving standard (typical) coin word problems
    - More complicated coin problems
    - Advanced coin problems
    - Solving coin problems mentally by grouping without using equations
    - Non-typical coin problems
    - Santa Claus helps solving coin problem
    - OVERVIEW of lessons on coin word problems
in this site.

You will find there the lessons for all levels - from introductory to advanced,
and for all methods used - from one equation to two equations and even without equations.

A convenient place to quickly observe these lessons from a  "bird flight height"  (a top view)  is the last lesson in the list.

Read them and become an expert in solution of coin problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Coin problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.